OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 60 miles per hour (mph), and the other car is traveling 50 mph. How long will it take before the cars are 220 miles apart?

OpenStudy (anonymous):

Since the cars are moving in opposite direction then their resultant speed = (60+50) mph = 110 mph Time = Distance/ speed Distance = 220 miles speed = 110 mph so time=distance/speed = 220/110 = 2 hours.

OpenStudy (foolaroundmath):

Suppose the time taken is 't' hours. Then distance traveled by first car = speed*time = 60t miles distance travelled by second car = 50t miles Separation between them = 60t + 50t = 110t (separation is added because they are moving away from each other) For separation to become 220 miles 110t = 220 => t = 2 hours